OpenStudy (anonymous):

The graphs of the function F (left, in blue) and G (right, in red) are below. Let P(x)=F(x)G(x) and Q(x)=F(x)/G(x). Answer the following questions. P'(1), Q'(1), P'(6), Q'(6)

OpenStudy (anonymous):

the first if F and the second is G

OpenStudy (anonymous):

@dumbcow @wio

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ash2326 @hartnn @Itsjustlaciee @thomaster

OpenStudy (watchmath):

Use product rule and quotient rule to compute \(P'(x)\) and \(Q'(x)\). To evaluate this at x=1,6 you need to read of F'(1),F'(6),G'(1),G'(6) from the given graph

OpenStudy (anonymous):

how do i find F'(1) and etc with only the graphs

OpenStudy (watchmath):

what F'(1) means geometrically?

OpenStudy (anonymous):

not sure

OpenStudy (watchmath):

is there any connection between derivative and tangent line?

OpenStudy (anonymous):

derivative is slope of tangent line

OpenStudy (dumbcow):

correct and lucky for you the graphs of F,G at points x=1, x=6 are straight lines so derivative is simply the slope of the line

OpenStudy (anonymous):

ok. so i find the slope at those points and then do what

OpenStudy (anonymous):

but is it a straight line for F

OpenStudy (dumbcow):

input those values into the product rule/quotient rule for F'(x), G'(x)

OpenStudy (dumbcow):

oh right F(1) is vertex of parabola, well imagine drawing a tangent line through that point....what is its slope?

OpenStudy (anonymous):

0

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

ok but how do i find F'(x) and G'(x)

OpenStudy (anonymous):

i dont know their equations

OpenStudy (dumbcow):

F'(x), G'(x) are just the slopes so we found slope at F(1) is 0 right...so F'(1) = 0

OpenStudy (anonymous):

ok and G'(1)=3

OpenStudy (dumbcow):

no look at slope of the line again

OpenStudy (anonymous):

rise/run=3/1

OpenStudy (dumbcow):

are you looking at line to right? when x=6 the line has slope of 3 G'(6) = 3

OpenStudy (anonymous):

oh sorry yeah its 1. i am doing both parts at the same time

OpenStudy (dumbcow):

ok then i get F'(6) = 1/4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would G'(1)=1?

OpenStudy (dumbcow):

yes you had it right

OpenStudy (anonymous):

ok thanks. i know how to get it now

OpenStudy (anonymous):

never mind i did something wrong

OpenStudy (anonymous):

F'(1)=0 F(1)=1 G'(1)=1 G(1)=2 F'(6)=1/4 F(6)=3 G'(6)=3 G(6)=3

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

this is what i got P'(1)=1 Q'(1)=1/4 P'(6)=39/4) Q'(6)=13/12